The magnetic field is 3T, the speed is 2m/s, the resistor is 70Ω and the width of the conducting bar is 7cm.

  1. What is the induced emf?
  2. What is the induced current?
  3. Is the current clockwise or counter clockwise?
Answer #1

given

magnetic filed B = 3 T

speed v = 2 m/s

resistancnce R = 70 ohm

A) induced emf V = Blv = 3*2*0.07 = 0.42 volts

B) current i = 0.42/70 = 0.006 A

c) counter clockwise